# Numerical Hopf–Lax formulae for Hamilton–Jacobi equations on unstructured geometries

### v1

@@ Abstract
We consider a scheme of Semi-Lagrangian (SL) type for the numerical solution of Hamilton–Jacobi (HJ) equation on unstructured triangular grids. As it is well known, SL schemes are not well suited for unstructured grids, due to the cost of the point location phase; this drawback is augmented by the need for repeated minimization. In this work, we propose a scheme that works only on the basis of node values and connectivity of the grid. In a first version, we obtain a monotone scheme; then, applying a quadratic refinement to the numerical solution, we improve accuracy at the price of some extra computational cost. The scheme can be applied to both time-dependent and stationary HJ equations; in the latter case, we also study the construction of a fast policy iteration solver. We perform a theoretical analysis of the two versions, and validate them with an extensive set of examples, both in the time-dependent and in the stationary case.
@@ 1 Introduction
The similarity between (4) and (1) is more apparent if we set $a=(x-y)/t$, and (4) is recast in the form
@@
which has been widely used for the SL discretization of (3), as well as of the Bellman equation of Dynamic Programming (see [6, 7] and the references therein). For (5), the SL discretization has the structure
@@
In this work, we will discretize again the representation formula (4) to obtain an approximation scheme for (3). However, in order to improve efficiency on the unstructured geometries to which this work is ultimately directed, we will drop the interpolation phase and perform minimization with a walk on the space grid, much in the spirit of [4]. We will then adapt this technique to the stationary case
$\lambda u+H(Du)=f(x)\quad x\in\mathbb{R}^{d},\$ (7)
@@
The outline of the paper is as follows. In Section 2 we will construct the scheme in the evolutive case and prove its convergence on the basis of the Barles–Souganidis theory [1], while Section 3 retraces the construction of the scheme for the stationary case (7). Last, we present a numerical validation in Section 6, and draw some conclusions in Section 7.
@@ 2 Construction of the scheme, time-dependent case
  • $H(\cdot)$ is $C^{2}$ with bounded second derivatives, strictly convex and strongly coercive;
  • $u_{0}$ is bounded and uniformly continuous.
@@
In the numerical approximation, we assume that the space grid covers the whole of $\mathbb{R}^{d}$. We will denote by $v_{j}^{n}$ the approximation of the solution at the node $x_{j}$ and time $t_{n}=n\Delta t$ (for $j\in\mathbb{N}$ and $n\geq 0$), and $V^{n}$ will stand for the infinite vector collecting all such values (we will later consider the equation on a bounded domain $\Omega$ with Dirichlet boundary conditions). We will avoid any assumption on both the regularity of the grid and the shape of the elements (although in the numerical examples they will be taken as triangles). We will also denote by $\Delta x$ the maximum diameter of the grid elements.
@@ 2.1 Restricting the Hopf–Lax formula to nodes
We start by rewriting the Hopf–Lax formula (4) on a single time step, from $t_{n}$ to $t_{n+1}$:
$u(x,t_{n+1})=\inf_{y\in\mathbb{R}^{d}}\left\{u(y,t_{n})+\Delta tH^{*}\left(\frac{x-y}{\Delta t}\right)\right\}.$ (8)
@@
$f(y)=u_{0}(y)+tH^{*}\left(\frac{x-y}{t}\right)$
is continuous and coercive, there exists a minimum point $\bar{y}$ for every $x\in\mathbb{R}^{d},\,t\in\mathbb{R}^{+}$. Moreover, given the basic assumptions above, the search for the minimum point can be confined to a compact set, as proved in [6].
@@
In previous works (see, e.g., [6]), the discretization of (8) has been carried out by taking $y\in\mathbb{R}^{d}$ and using a descent-type minimization algorithm, with an interpolated value for $u(y,t_{n})$. In this case, at the price of some drop in accuracy, we avoid both phases of interpolation and iterative minimization, which would result in a high computational cost. We remark that this use of the Hopf–Lax formula has some precedent (for example, [3]), although, to the Authors’ knowledge, no attempt to optimize the algorithm for unstructured grids has been made so far.
Convergence of this scheme will be analysed now in the framework of Barles–Souganidis theory [1]. As it is well-known, this convergence theory requires the scheme to be monotone, $L^{\infty}$-stable and consistent. In our case, the scheme is invariant for the addition of constants, since, for any $c\in\mathbb{R}$,
@@
Assume now that $V$ and $W$ are two vectors of nodal values such that $V\geq W$ element by element. We want to prove that
@@
Let now $u$ be a smooth solution of (3). Consistency of the scheme requires that for all $j$, as $\Delta x,\Delta t\to 0$,
@@
$\mathcal{F}_{j}(y)=u(y,t_{n})+\Delta tH^{*}\left(\frac{x_{j}-y}{\Delta t}\right)$ (14)
and its minimum restricted to a grid with step $\Delta x$; for simplicity, we will carry out the computation in one space dimension.
@@
The first term in the right-hand side is bounded for a smooth solution, while the second, given the boundedness of ${H^{*}}^{\prime\prime}$, is of order $1/\Delta t$. Then, the difference between the minimum of $\mathcal{F}_{j}$ and its minimum value at the nodes is $O(\|{H^{*}}^{\prime\prime}\|_{\infty}\Delta x^{2})$, and the consistency error turns out to be
@@
Therefore, the scheme is consistent under an inverse CFL condition $\Delta x=o(\Delta t)$. For example, taking the relationship $\Delta t\sim\Delta x^{1/2}$, the scheme turns out to be first-order w.r.t. $\Delta x$.
@@
Let the basic assumptions hold, and assume in addition that $\Delta x=o(\Delta t)$. Then, the numerical solution $V^{n}$ defined by (9) converges to the solution $u(x,t_{n})$ of (3) locally uniformly on $\mathbb{R}^{d}\times[0,T]$ as $\Delta t,\Delta x\to 0$ for any $T>0$.
Remark 1
@@
In case weak Dirichlet conditions (see [7]) are enforced in (3), assume that the boundary datum $b_{j}=b(x_{j})$ is assigned at a boundary node $x_{j}$ for a given function $b$. Then, the update formula (9) for $v_{j}$ is replaced by
@@
$v_{j}^{n+1}=\min\left\{b_{j},\min_{k}\left\{v_{k}^{n}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}\right\}.$
In the dynamic programming setting, this amounts to set a stopping cost on the boundary of the domain. The above convergence analysis can be replicated with minor modifications.

### v2

@@ Abstract
We consider a scheme of Semi-Lagrangian (SL) type for the numerical solution of Hamilton–Jacobi (HJ) equations on unstructured triangular grids. As it is well known, SL schemes are not well suited for unstructured grids, due to the cost of the point location phase; this drawback is augmented by the need for repeated minimization. In this work, we consider an existing, monotone version of the scheme, that works only on the basis of node values, and adapt the algorithm to the case of an unstructured grid, using the connectivity information. Then, applying a quadratic refinement to the numerical solution, we improve accuracy at the price of some extra computational cost. The scheme can be applied to both time-dependent and stationary HJ equations; in the latter case, we also study the construction of a fast policy iteration solver and of a parallel version. We perform a theoretical analysis of the two versions, and validate them with an extensive set of examples, both in the time-dependent and in the stationary case.
@@ 1 Introduction
The similarity between (4) and (1) is more apparent if we set a=(xy)/ta=(x-y)/t, and (4) is recast in the form
@@
which has been widely used for the SL discretization of (3), as well as of the Bellman equation of Dynamic Programming (see [8, 9] and the references therein).
The numerical discretization of (3) has been extensively studied; a general review on the topic is provided in [10]. However, a relatively small part of this literature is devoted to the construction of efficient schemes on complex geometries, which require unstructured space grids. Among the relevant works in this direction, we quote in particular the researches on monotone, ENO–WENO and Discontinuous Galerkin schemes (see the review [19]) and those using representation formulae of Hopf–Lax or Dynamic Programming type [4, 13, 18].
This latter technique has the advantage of allowing the use of large time steps, and has therefore been used to construct schemes of semi-Lagrangian type. For (5), the typical SL discretization has the structure
@@
In this work, we will discretize again the representation formula (4) to obtain an approximation scheme for (3), but drop the interpolation phase and restrict the computation to nodes. We remark that this use of the Hopf–Lax formula has various precedents (for example, [5, 12]; however, our interest here is to improve efficiency on unstructured grids to which this work is ultimately directed. To this end, the basic idea will be to perform minimization with a walk on the space grid, much in the spirit of [6].We will then adapt this technique to problems with Dirichlet boundary conditions and to the stationary case
$\lambda u+H(Du)=f(x),$ (7)
@@
The outline of the paper is as follows. In Section 2 we will construct the scheme in the evolutive case and prove its convergence on the basis of the Barles–Souganidis theory [2], while Section 3 retraces the construction of the scheme for the stationary case (7). Sections 45 treat respectively the unstructured implementation and a technique of quadratic refinement to improve accuracy of the scheme. Last, we present a numerical validation in Section 6, and draw some conclusions in Section 7.
@@ 2 Construction of the scheme, time-dependent case
  • $H(\cdot)$ is $C^{2}$ with bounded second derivatives, so that
    $\max_{p}|D^{2}H(p)|_{2}\leq M_{H};$
    moreover, it is strictly convex and strongly coercive;
  • $u_{0}$ is bounded and Lipschitz continuous.
@@
In the numerical approximation, we start by assuming that the space grid covers the whole of $\mathbb{R}^{d}$. We will denote by $v_{j}^{n}\quad(j,n\in\mathbb{N})$ the approximation of the solution at the node $x_{j}$ and time $t_{n}=n\Delta t$ (for $j\in\mathbb{N}$ and $n\geq 0$), and $V^{n}$ will stand for the infinite vector collecting all such values. We avoid any assumption on both the regularity of the grid and the shape of the elements (although in the numerical examples they will be taken as triangles). We will also denote by $\Delta x$ the maximum diameter of the grid elements.
@@ 2.1 Restricting the Hopf–Lax formula to nodes
We start by rewriting the Hopf–Lax formula (4) at a node $x_{j}$ and on a single time step, from $t_{n}$ to $t_{n+1}$:
$u(x_{j},t_{n+1})=\inf_{y\in\mathbb{R}^{d}}\left\{u(y,t_{n})+\Delta tH^{*}\left(\frac{x_{j}-y}{\Delta t}\right)\right\}.$ (8)
@@
$\mathcal{F}_{j}^{0}(y,t)=u_{0}(y)+tH^{*}\left(\frac{x_{j}-y}{t}\right)$
is continuous and coercive, there exists a minimum point $\bar{y}$ for every $x\in\mathbb{R}^{d},\,t\in\mathbb{R}^{+}$. Moreover, given the basic assumptions above, the search for the minimum point can be restricted to a compact set, as proved in [8].
@@
In previous works (see, e.g., [8]), the discretization of (8) has been carried out by taking $y\in\mathbb{R}^{d}$ and using a descent-type minimization algorithm, with an interpolated value for $u(y,t_{n})$. In this case, at the price of some drop in accuracy, we avoid both phases of interpolation and iterative minimization, which would result in a high computational cost.
For the moment, we will neglect boundary conditions, and analyse convergence of the scheme assuming an infinite grid covering the whole of $\mathbb{R}^{d}$, whereas in a second phase we will examine the case of a bounded domain. Here, we will work in the framework of Barles–Souganidis theory [2]. As it is well-known, this convergence theory requires the scheme to be monotone, $L^{\infty}$-stable and consistent. In our case, the scheme is invariant for the addition of constants, since, for any $c\in\mathbb{R}$,
@@
Monotonicity of the scheme has already been proved elsewhere, but we include the proof for completeness. Assume that $V$ and $W$ are two vectors of nodal values such that $V\geq W$ element by element. We want to prove that
@@
Let now $u$ be a smooth solution of (3). In its usual sense, consistency of the scheme requires that for all $j$, as $\Delta x,\Delta t\to 0$,
@@
$\mathcal{F}_{j}^{n}(y,\Delta t)=u(y,t_{n})+\Delta tH^{*}\left(\frac{x_{j}-y}{\Delta t}\right)$ (14)
and its minimum restricted to a grid with step $\Delta x$; for simplicity, we will carry out the computation in one space dimension and use the shorthand notation $\mathcal{F}_{j}(y)$ for (14).
@@
The first term in the right-hand side is bounded for a smooth solution, while the second, given the boundedness of $H^{*}$$^{\prime\prime}$, is of order $1/\Delta t$. Then, the difference between the minimum of $\mathcal{F}_{j}$ and its minimum value at the nodes is $O(\|{H^{*}}^{\prime\prime}\|_{\infty}\Delta x^{2})$, and the consistency error reads
@@
Therefore, the scheme is consistent under an inverse CFL condition $\Delta x=o(\Delta t)$ (this latter consistency condition also appears in [5, 12]). For example, taking the relationship $\Delta t\sim\Delta x^{1/2}$, the scheme turns out to be first-order w.r.t. $\Delta x$.
Remark 1
The previous study of the local truncation error aims at characterizing the accuracy of the scheme, while [2] adopts a weaker concept of consistency. More precisely, the Barles–Souganidis theorem would require that, for any function $\phi\in C^{\infty}(\mathbb{R}^{d}\times[0,T])$,
$\lim_{\Delta t\to 0}\frac{\displaystyle\phi(x_{j},t_{n+1})-\min_{k}\left\{\phi(x_{k},t_{n})+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}}{\Delta t}=\phi_{t}(x_{j},t_{n})+H(D\phi(x_{j},t_{n})).$ (15)
However, we show that (15) also follows from the previous arguments. In fact, writing $H$ as the Legendre transform of $H^{*}$, we have
$\phi_{t}(x_{j},t_{n})+H(D\phi(x_{j},t_{n}))$ $=$ $\phi_{t}(x_{j},t_{n})+\max_{q\in\mathbb{R}^{d}}\left\{q\cdot D\phi(x_{j},t_{n})-H^{*}(q)\right\}$
$=$ $\frac{\phi(x_{j},t_{n+1})-\phi(x_{j},t_{n})}{\Delta t}$
$+\max_{y}\left\{\frac{\phi(x_{j},t_{n})-\phi(y,t_{n})}{\Delta t}-H^{*}\left(\frac{x_{j}-y}{\Delta t}\right)\right\}+O(\Delta t),$
where we have and approximated $\phi_{t}$ via the time incremental ratio, the directional derivative $q\cdot D\phi$ with the incremental ratio with step $\Delta t$ in the direction
$q=\frac{x_{j}-y}{\Delta t},$
and collected all the error terms in a $O(\Delta t)$. Simplifying now the terms $\pm\phi(x_{j},t_{n})/\Delta t$ and rewriting the $\max$ as a $\min$, we finally get
$\phi_{t}(x_{j},t_{n})+H(D\phi(x_{j},t_{n}))=\frac{\displaystyle\phi(x_{j},t_{n+1})-\min_{y}\left\{\phi(y,t_{n})+\Delta tH^{*}\left(\frac{x_{j}-y}{\Delta t}\right)\right\}}{\Delta t}+O(\Delta t).$ (16)
Using (16) in the right-hand side of (15), consistency in the sense of [2] finally comes down to comparing $\mathcal{F}_{j}^{n}(y,\Delta t)$ with $\mathcal{F}_{j}^{n}(x_{k},\Delta t)$, and may be obtained by the arguments already used above.
@@
Let the basic assumptions hold, and assume, in addition, that $\Delta x=o(\Delta t)$. Then, the numerical solution $V^{n}$ defined by (9) converges to the solution $u(x,t_{n})$ of (3) locally uniformly on $\mathbb{R}^{d}\times[0,T]$ as $\Delta t,\Delta x\to 0$ for any $T>0$.
2.2 Error estimates
In order to obtain explicit error estimates for the numerical solution $V^{n}$, we recall that the exact solution $u$ of (3) is Lipschitz continuous (with, say, a Lipschitz constant $L_{u}$), and that the speed of propagation $(x-y)/\Delta t$ achieving the minimum in (4) is bounded, i.e., that we can restrict the argument of $H^{*}$ to a bounded set, in which $H^{*}$ is also Lipschitz continuous, with a Lipschitz constant $L_{H^{*}}$. Then, comparing (8) and (9) and denoting respectively by $y_{j}$ and $k_{j}$ the two minimizers, we have a first one-sided bound of the form
$u(x_{j},t_{n+1})-v_{j}^{n+1}$ $=$ $u(y_{j},t_{n})+\Delta tH^{*}\left(\frac{x_{j}-y_{j}}{\Delta t}\right)-v_{k_{j}}^{n}-\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)$ (17)
$\leq$ $u(x_{k_{j}},t_{n})+\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)-v_{k_{j}}^{n}-\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)$
$\leq$ $u(x_{k_{j}},t_{n})-v_{k_{j}}^{n},$
where we have used the fact that $x_{k_{j}}$ is not in general a minimizer in (8). Iterating back (17) and taking into account that $v_{j}^{0}=u_{0}(x_{j})$, we obtain
$u(x_{j},t_{n+1})-v_{j}^{n+1}\leq 0.$
As for the reverse estimate, we denote by $x_{k_{j}^{*}}$ the grid node closest to $y_{j}$, so that $|x_{k_{j}^{*}}-y_{j}|\leq\Delta x$. Then, we get:
$v_{j}^{n+1}-u(x_{j},t_{n+1})$ $=$ $v_{k_{j}}^{n}+\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)-u(y_{j},t_{n})-\Delta tH^{*}\left(\frac{x_{j}-y_{j}}{\Delta t}\right)$ (18)
$\leq$ $v_{k_{j}^{*}}^{n}+\Delta tH^{*}\left(\frac{x_{j}-x_{k_{j}^{*}}}{\Delta t}\right)-u(y_{j},t_{n})-\Delta tH^{*}\left(\frac{x_{j}-y_{j}}{\Delta t}\right)$
$=$ $\left(v_{k_{j}^{*}}^{n}-u(x_{k_{j}^{*}},t_{n})\right)+\left(u(x_{k_{j}^{*}},t_{n})-u(y_{j},t_{n})\right)$
$+\Delta t\left|H^{*}\left(\frac{x_{j}-x_{k_{j}^{*}}}{\Delta t}\right)-H^{*}\left(\frac{x_{j}-y_{j}}{\Delta t}\right)\right|$
$=$ $v_{k_{j}^{*}}^{n}-u(x_{k_{j}^{*}},t_{n})+L_{u}\Delta x+\Delta tL_{H^{*}}\frac{\Delta x}{\Delta t}.$
Finally, denoting by $U^{n}$ the vector of node values of the exact solution at $t_{n}$, iterating back (18) for $t_{n}\leq T$ and including (17), we obtain the estimate:
$0\leq V^{n}-U^{n}\leq T(L_{u}+L_{H^{*}})\frac{\Delta x}{\Delta t},$
where the inequalities are clearly understood component-wise.
2.3 Boundary conditions
@@
In case Dirichlet conditions are enforced in (3), we assume that the HJ equation is posed on a bounded open set $\Omega\subset\mathbb{R}^{d}$, and that the boundary datum $b(x)$ is assigned at any boundary point $x\in\partial\Omega$, for a given continuous function $b$. Then, the HJ equation takes the form
$\begin{cases}u_{t}+H(Du)=0&(x,t)\in\Omega\times\mathbb{R}^{+},\\ u(x)=b(x)&x\in\partial\Omega,\\ u(x,0)=u_{0}(x),&x\in\Omega.\end{cases}$ (19)
We recall that the Dirichlet condition in (19) should be understood in the weak sense (see [9] for details). In this case, up to a shift and change of sign of the time variable, the solution of (19) may be characterized, via dynamic programming arguments (see [1]), as the value function of the optimal control problem
$\min_{\alpha(\cdot)}\int_{t}^{T\wedge\tau_{\alpha}(x,t)}H^{*}(\alpha(s))ds+u_{0}(X_{\alpha}(T\wedge\tau_{\alpha}(x,t))),$ (20)
where $T>0$, $t\in[0,T]$, $\alpha\in L^{1}(\mathbb{R}^{+},\mathbb{R}^{d})$ and $X_{\alpha}(\cdot)$ satisfies
$\begin{cases}\dot{X}_{\alpha}(s)=\alpha(s)&s\geq t,\\ X_{\alpha}(t)=x\in\overline{\Omega}.\end{cases}$ (21)
Here, $\tau_{\alpha}(x,t)$ is the exit time associated to the initial condition $(x,t)$ and the control $\alpha$, defined as
$\tau_{\alpha}(x,t):=\inf\{s\geq t:X_{\alpha}(s)\in\partial\Omega\},$ (22)
and we use the notation $a\wedge b=\min\{a,b\}$. We also assume that the final cost $u_{0}$ is compatible with the boundary stopping cost $b$, i.e., that $u_{0}(x)=b(x)$ for $x\in\partial\Omega$.
Then, setting $\alpha(s)\equiv(x_{j}-y)/(\Delta t\wedge\tau)$, for $x\in\overline{\Omega}$ the Hopf–Lax formula in the form (8) is replaced by
$u(x_{j},t_{n+1})=\min_{y\in\overline{\Omega}\atop\tau\in(0,\Delta t]}\mathcal{F}_{j}^{n}(y,\tau)$ (23)
where
[…]
and the fully discrete scheme is finally obtained as
$v_{j}^{n+1}=\min_{x_{k}\in\overline{\Omega}\atop\tau\in(0,\Delta t]}F_{j}^{n}(x_{k},\tau),$ (25)
in which $F_{j}^{n}$ is the fully discrete version of $\mathcal{F}_{j}^{n}$, defined in turn as
[…]
The discretization (25)–(26) differs from (9) only when $y=x_{k}\in\partial\Omega$ – in this case, a minimization with respect to the exit time is also required. Therefore, for each $x_{k}$ on the boundary, we must compute
$\min_{\tau\leq\Delta t}\mathcal{F}_{j}^{n}(x_{k},\tau)=\min_{\tau\leq\Delta t}\left\{b(x_{k})+\tau H^{*}\left(\frac{x_{j}-x_{k}}{\tau}\right)\right\}.$
The term to be minimized is smooth and strictly convex, therefore its minimum is attained at the minimum between its only stationary point $\tau^{*}$ and $\Delta t$ (we will denote this optimal value as $\tau_{jk}$). We have:
$\frac{d}{d\tau}\left\{v_{k}^{n}+\tau H^{*}\left(\frac{x_{j}-x_{k}}{\tau}\right)\right\}=H^{*}\left(\frac{x_{j}-x_{k}}{\tau}\right)-\frac{x_{j}-x_{k}}{\tau}\cdot\nabla H^{*}\left(\frac{x_{j}-x_{k}}{\tau}\right),$ (27)
which allows to find the stationary point, either explicitly or numerically, as soon as the analytic form of $H^{*}$ is known, via a restriction of $H^{*}$ to the direction $x_{j}-x_{k}$. In particular, in the numerical tests we will use the hamiltonian
$H(p)=a_{0}+\frac{1}{2}p^{2},$
whose Legendre transform is
$H^{*}(q)=\frac{1}{2}q^{2}-a_{0}.$
Using this form of $H^{*}$, the stationary point in (27) is computed, after some algebra, by solving the stationarity condition
$a_{0}-\frac{1}{2}\frac{(x_{j}-x_{k})^{2}}{\tau^{2}}=0,$
which has no solution if $a_{0}>0$ (in which case the optimal stopping time is $\Delta t$), and otherwise has the solution
$\tau^{*}=\frac{|x_{j}-x_{k}|}{\sqrt{-2a_{0}}}.$
The time step to be used with the node $x_{k}$ is therefore
@@
$\tau_{jk}=\begin{cases}\Delta t&\text{ if }a_{0}>0,\\ \min\left\{\Delta t,\frac{|x_{j}-x_{k}|}{\sqrt{-2a_{0}}}\right\}&\text{ otherwise.}\end{cases}$